i need help: what is the solution set of the equation x/5 + x/2=14?: i need help: what is the solution set of the equation x/5 + x/2=14?
Answer by stanbon(19743)

what is the solution set of the equation x/5 + x/2=14?
-----------
Multiply thru by 70 to get:
14x + 35x = 5
49x = 5
x = 5/49
============
Cheers,
Stan H.